OpenStudy (anonymous):

ok, so i get what an arithmetic thingy is now. but how do you find one with fractions?! State whether the following sequence is arithmetic; and if so, state the common difference. 1/9, 1/3, 5/9, 7/9, ... 2/9 -2/9 not arithmetic 1/3

OpenStudy (anonymous):

1/3-1/9 2/9

OpenStudy (anonymous):

5/9-1/3 2/9

OpenStudy (anonymous):

what middle?

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

1/3 = 3/9 so 1/9, 3/9, 5/9, 7/9.. arithmatic series.. diff = 2/9

OpenStudy (anonymous):

SINCE ITS AN AP ,the difference between every two terms of the sequence is 2/9

OpenStudy (anonymous):

3/9 - 1/9 = 5/9 - 3/9 = 7/9 - 5/9 = 2/9 It is an Arithmetic Progression with Common Difference of 2/9..